# Kinematify: Open-Vocabulary Synthesis of High-DoF Articulated Objects

### v1

@@ Abstract
A deep understanding of kinematic structures and movable components is essential for enabling robots to manipulate objects and model their own articulated forms. Such understanding is captured through articulated objects, which are essential for tasks such as physical simulation, motion planning, and policy learning. However, creating these models, particularly for complex systems like robots or objects with high degrees of freedom (DoF), remains a significant challenge. Existing methods typically rely on motion sequences or strong assumptions from hand-curated datasets, which hinders scalability. In this paper, we introduce Kinematify, an automated framework that synthesizes articulated objects directly from arbitrary RGB images or text prompts. Our method addresses two core challenges: (i) inferring kinematic topologies for high-DoF objects and (ii) estimating joint parameters from static geometry. To achieve this, we combine MCTS search for structural inference with geometry-driven optimization for joint reasoning, producing physically consistent and functionally valid descriptions. We evaluate Kinematify on diverse inputs from both synthetic and real-world environments, demonstrating improvements in registration and kinematic topology accuracy over prior work.
@@ I Introduction
Enabling robots to effectively interact with and manipulate objects, as well as to model their own articulated structures for self-perception and adaptation, requires a deep understanding of kinematic topologies and joint parameters. Articulated robot descriptions capture this understanding by encoding geometry, kinematic dependencies, and dynamic constraints in standard formats like the Unified Robot Description Format (URDF) [1]. These descriptions are essential for core robotic tasks such as manipulation, locomotion, and policy learning. However, customizing such descriptions for articulated objects remains a significant challenge, demanding substantial manual effort, especially for high degrees of freedom (DoF) systems like humanoids, quadrupeds, and arms. This difficulty arises from the labor intensive processes of part-aware 3D modeling, resolving intricate kinematic dependencies, and inferring precise joint parameters [2].
@@
While recent advances in part-aware 3D generation [3, 4, 5, 6, 7] now enable the on-demand creation of high-quality segmented meshes from text or image prompts, the bottleneck of kinematics inference remains. This challenge has driven robotics researchers to explore high-DoF articulated objects generation approaches.
@@
Prior work has followed two main directions. Geometry-first approaches infer parts and joints from dense 4D sequences or multi-scan data, which achieve high fidelity but rely on controlled capture settings [8, 9]. Program-synthesis pipelines, in contrast, predict executable descriptions directly from visual inputs [10, 11, 12]. While effective, these systems mainly target everyday objects such as laptops, bottles, and drawers, which typically contain only a few moving parts and relatively simple kinematic dependencies. In the context of robot self-modelling, related work such as AutoURDF [13] extends this idea to robots, recovering topology and joint types from point-cloud sequences. However, it presumes motion data and is largely limited to serial-chain structures, whereas high-DoF objects often exhibit multi-branched linkages.
@@
We introduce Kinematify, a pipeline that synthesizes articulated robot descriptions from RGB images or texts. Our method first generates the segmented mesh with a 3D foundation model, then constructs a contact graph and infers a kinematic tree using an MCTS [14, 15, 16] objective that balances hierarchy and structural regularity. Subsequently, it estimates joint parameters via DW-CAVL, a novel Distance-Weighted Contact-Aware Virtual Linkage optimization approach. This approach preserves near-contact regions while penalizing collisions under virtual motion. The resulting description is exported to URDF and is readily convertible to formats like MJCF [17] or USD [18]. Our contributions are:
@@
  • An automated pipeline for articulated object synthesis. Kinematify generates physics-aware, high-DoF articulated objects directly from arbitrary RGB images or text prompts, without requiring motion data or pre-defined articulation priors.
@@
  • MCTS-based kinematic tree inference. We propose a search objective that encodes structural priors like hierarchy and regularity to resolve ambiguous attachments for complex, high-DoF articulated objects.
@@
  • A geometry-driven joint parameter estimation method. Our DW-CAVL algorithm accurately infers revolute and prismatic joint parameters from static geometry alone by optimizing an SDF-based, contact-aware objective under virtual motions.
@@ II Related Work
A significant body of work focuses on reconstructing the kinematic structure of everyday objects from visual data. The most common paradigm leverages motion to reveal articulation [9, 8, 19, 20, 13]. By observing an object over time, these methods can directly infer which parts move together and identify the axes of rotation or translation. For example, MultiBodySync [9] registers multiple 3D scans of an object in different states, using spectral synchronization to jointly solve for part segmentation and motion. Similarly, ReArt [8] fits a rearticulable model to a 4D point cloud sequence by jointly optimizing segmentation, topology, and joint parameters. These methods achieve high fidelity but depend critically on the availability of multi-view or temporal data, which requires a controlled capture setup. Another trend frames articulation modelling as a program synthesis problem [12, 10, 11, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. URDFormer [12] trains a transformer to predict a URDF from a single image, relying on a large-scale synthetic dataset of image-URDF pairs. Real2Code [10] and Articulate-Anything [11] leverage large language models (LLMs) to generate code-based representations of articulated objects, with the latter using a reinforcement learning loop to refine the model through simulation feedback. While powerful in their open-vocabulary capabilities, these methods often produce functionally plausible rather than geometrically precise models and struggle with the complex topologies common in high-DoF objects.
@@ II-B Robot Self-Modeling
In contrast to robot self-modelling methods that require an observation of the robot, Kinematify operates in a zero-shot physical context, inferring the kinematics topology from the image or text alone. These open vocabulary settings make our approach applicable much earlier in the design pipeline, before a physical prototype is built or motion data is available. To the best of our knowledge, Kinematify is the first method to construct high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or texts.
@@ III Kinematify
To overcome the limitations of prior self-modelling methods, we introduce Kinematify, a zero-shot framework that constructs high-DoF, kinematics-aware articulated objects directly from unlabelled RGB images or textual descriptions.
@@ III-A Preliminaries
Assemblies and parts. An assembly $\mathcal{A}$ is a finite set of parts $P=\{P_{i}\}_{i=1}^{N}$. Part $P_{i}$ has a triangulated surface mesh $M_{i}=(V_{i},F_{i})$ with vertices $V_{i}\subset\mathbb{R}^{3}$ and triangular faces $F_{i}\subset\{1,\dots,|V_{i}|\}^{3}$. Each part stores a world transform $T_{i}\in\mathrm{SE}(3)$, an intrinsic rotation $R_{i}\in\mathrm{SO}(3)$ used for alignment, a centroid $c_{i}\in\mathbb{R}^{3}$, a robust volume $v_{i}>0$, and axis-aligned bounding-box extents $e_{i}\in\mathbb{R}^{3}_{>0}$.
Graphs and kinematic tree. We build an undirected connection graph $G=(V,E)$ with $V\leftrightarrow P$; an edge indicates geometric contact. A directed kinematic tree $T=(V,E_{T})$, rooted at the base link $b\in V$, orients $G$ and annotates joints.
@@ III-B Part-Aware 3D Representations
We generate 3D part-level 3D meshes with a 3D foundation model [3] from the input RGB image, and discard meshes with too few vertices or a degenerate spatial spread. For each prospective parent part, we train a continuous SDF [39, 40, 41] $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$ on (i) noisy surface points, (ii) near-surface offsets, and (iii) far samples in a bounding AABB.
@@ III-C Kinematic Topology Inference
We select the base link as any
$b\in\arg\max_{i\in V(G)}\deg_{G}(i),$
where $\deg_{G}(i)$ is the undirected degree in $G$. Starting from $b$, we run BFS on $G$ as a warm start for the MCTS search. During BFS, when a new neighbour $v$ is first reached from an already visited node $u$, we define $u$ as the parent and $v$ as the child, orient the edge as $u\!\to\!v$, and set its origin
$o_{uv}=c_{v}-c_{u}.$ (1)
Any edge whose addition would create a cycle is not inserted into $T$ and is recorded as broken. If $G$ is disconnected, each remaining component is attached to the current tree by adding a virtual edge from its nearest neighbour in the tree, provided the Euclidean distance is at most $d_{\max}$.
@@ III-C1 State, actions, constraints
Figure 1: Pipeline of Kinematify for recovering articulated robots from a single RGB image. Step 1: A 3D foundation model generates a segmented mesh of the robot. Step 2: A contact graph is constructed over mesh parts, capturing candidate relations between components. Step 3: MCTS is applied to infer the kinematic tree, resolving ambiguous attachments by leveraging structural priors such as hierarchy and symmetry. Step 4: The DW-CAVL optimization refines joint parameters by preserving near-contact geometry. Bottom row: examples of inferred revolute and prismatic joints with optimized axes.
@@ III-C3 Reward
$R_{\mathrm{struct}}$.
This term penalises irregular depths, excessive out-degree, and long trees:
$R_{\mathrm{struct}}=\frac{1}{1+\overline{d^{2}}+\overline{(\deg^{+}-3)^{2}}+\lambda\,|E(\tilde{T})|}$, (2)
where $\lambda>0$ is a hyperparameter.
$R_{\mathrm{static}}$.
@@
where $\mathrm{ch}(i)$ is the set of children of $i$ in $\tilde{T}$. Let $\hat{\mathbf{z}}^{-}=[0,\,0,\,-1]^{\top}$, which represents downward unit gravity and $g>0$ the gravitational constant. The total gravitational torque is
@@ III-D2 DW-CAVL objective
The consistency term penalises separation near contact after motion,
@@ IV-D Ablation study
Figure 5: Kinematify generates URDFs for both the Fetch robot and the drawer, enabling a demonstration of the robot opening the drawer in Isaac Sim and on hardware, with the same models usable for online planning in ROS and MoveIt.
@@ IV-E Real-world robot manipulation
V Limitations
Two limitations remain. First, our approach assumes that contact regions are reliable articulation cues and that the contact graph is connected. Spurious contacts can mislead the search and leave parts unattached. Second, we restrict the recovered kinematics to a tree structure and thus do not support closed-loop mechanisms, such as four-bar linkages. Future work includes contact reliability estimation and outlier rejection, and extending the solver with loop-closure constraints to handle kinematic loops.
@@ VI Conclusion
We presented Kinematify, an automated pipeline that synthesizes articulated object and robot descriptions directly from RGB images or text prompts. Across both synthetic and real-world inputs, Kinematify improves registration quality and kinematic topology over prior work. We view Kinematify as a step toward open-vocabulary synthesis of high-DoF articulated structures.

### v2

@@ Abstract
A deep understanding of kinematic structures and movable components is essential for enabling robots to manipulate objects and model their own articulated forms. Such understanding is captured through articulated objects, which are essential for tasks such as physical simulation, motion planning, and policy learning. However, creating these models, particularly for objects with high degrees of freedom (DoF), remains a significant challenge. Existing methods typically rely on motion sequences or strong assumptions from hand-curated datasets, which hinders scalability. In this paper, we introduce Kinematify, an automated framework that synthesizes articulated objects directly from arbitrary RGB images or textual descriptions. Our method addresses two core challenges: (i) inferring kinematic topologies for high-DoF objects and (ii) estimating joint parameters from static geometry. To achieve this, we combine MCTS search for structural inference with geometry-driven optimization for joint reasoning, producing physically consistent and functionally valid descriptions. We evaluate Kinematify on diverse inputs from both synthetic and real-world environments, demonstrating improvements in registration and kinematic topology accuracy over prior work. https://sites.google.com/deemos.com/kinematify
@@ I Introduction
Enabling robots to effectively interact with objects, as well as to model their own articulated structures for self-perception and adaptation, requires an accurate understanding of kinematic topologies and joint parameters. Articulated robot descriptions capture this understanding by encoding geometry, kinematic dependencies, and dynamic constraints in standard formats like the Unified Robot Description Format (URDF) [1]. These descriptions are essential for robotic tasks such as manipulation, locomotion, and policy learning. However, customizing such descriptions for articulated objects remains a significant challenge, demanding substantial manual effort, especially for high degrees of freedom (DoF) systems like humanoids, quadrupeds, and arms. This difficulty arises from the labor intensive processes of part-aware 3D modeling, resolving intricate kinematic dependencies, and inferring precise joint parameters [2].
@@
While recent advances in part-aware 3D generation [3, 4, 5, 6, 7] now enable the on-demand creation of high-quality segmented meshes from RGB images or textual descriptions, the bottleneck of kinematics inference remains. This challenge has driven robotics researchers to explore high-DoF articulated objects generation approaches.
@@
Prior work has followed two main directions. Geometry-first approaches infer parts and joints from dense 4D sequences or multi-scan data, which achieve high fidelity but rely on controlled capture settings [8, 9]. Program-synthesis pipelines, in contrast, predict executable descriptions directly from visual inputs [10, 11, 12]. While effective, these systems mainly target everyday objects such as laptops, bottles, and drawers, which typically contain only a few moving parts and relatively simple kinematic dependencies. In the context of self-modeling, related work such as AutoURDF [13] extends this idea to robots, recovering topology and joint types from point-cloud sequences. However, it presumes motion data and is largely limited to serial-chain structures, whereas high-DoF objects often exhibit multi-branched linkages.
@@
To address these challenges, we introduce Kinematify, a framework that generates articulated 3D objects from RGB images or texts. Kinematify generates the segmented mesh with a part-aware 3D foundation model, such as Rodin [3], then infers the kinematic tree using an MCTS [14, 15, 16] objective that balances hierarchy and structural regularity. Subsequently, it estimates joint parameters via DW-CAVL, a novel Distance-Weighted Contact-Aware Virtual Linkage optimization approach. This approach preserves near-contact regions while penalizing collisions under virtual motion. The resulting description is exported to URDF and is readily convertible to formats like MJCF [17] or USD [18]. Our contributions are:
@@
  • An open-vocabulary articulated object generation framework. Kinematify generates physics-aware articulated objects directly from arbitrary RGB images or textual descriptions, without requiring motion data, training, or pre-defined articulation priors.
@@
  • A MCTS-based kinematic tree inference approach. We propose a search objective that encodes structural priors like hierarchy and regularity to resolve ambiguous attachments for complex, high-DoF articulated objects with multiple branches.
@@
  • A SDF-driven joint parameter estimation approach. The DW-CAVL algorithm accurately infers revolute and prismatic joint parameters from static geometry by optimizing an SDF-based, contact-aware objective under virtual motions.
@@ II Related Work
A significant body of work focuses on reconstructing the kinematic structure of everyday objects from visual data. The most common paradigm leverages motion to reveal articulation [9, 8, 19, 20, 13]. By observing an object over time, these methods can directly infer which parts move together and identify the axes of rotation or translation. For example, MultiBodySync [9] registers multiple 3D scans of an object in different states, using spectral synchronization to jointly solve for part segmentation and motion. Similarly, ReArt [8] fits a rearticulable model to a 4D point cloud sequence by jointly optimizing segmentation, topology, and joint parameters. These methods achieve high fidelity but depend critically on the availability of multi-view or temporal data, which requires a controlled capture setup. Another trend frames articulation modelling as a program synthesis problem [12, 10, 11, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. URDFormer [12] trains a transformer to predict a URDF from a single image, relying on a large-scale synthetic dataset of image-URDF pairs. Real2Code [10] and Articulate-Anything [11] leverage large language models to generate code-based representations of articulated objects, with the latter using a reinforcement learning loop to refine the model through simulation feedback. While powerful in their open-vocabulary capabilities, these methods often produce functionally plausible rather than geometrically precise models and struggle with the multi-branch kinematics, which are common in high-DoF objects.
@@ III Kinematify
We introduce Kinematify, a framework that generates kinematics-aware articulated objects directly from RGB images or textual descriptions in a zero-shot context.
@@ III-A Preliminaries
Assemblies and parts. An assembly $\mathcal{A}$ is a set of parts $P=\{P_{i}\}_{i=1}^{N}$. Part $P_{i}$ has a triangulated surface mesh $M_{i}=(V_{i},F_{i})$ with vertices $V_{i}\subset\mathbb{R}^{3}$ and triangular faces $F_{i}\subset\{1,\dots,|V_{i}|\}^{3}$. Each part stores a world transform $T_{i}\in\mathrm{SE}(3)$, an intrinsic rotation $R_{i}\in\mathrm{SO}(3)$ used for alignment, a centroid $c_{i}\in\mathbb{R}^{3}$, a robust volume $v_{i}>0$, and axis-aligned bounding-box extents $e_{i}\in\mathbb{R}^{3}_{>0}$.
Graphs and kinematic tree. We build an undirected connection graph $G=(V,E)$ with $V\leftrightarrow P$, where an edge indicates geometric contact. A directed kinematic tree $T=(V,E_{T})$, rooted at the base link $b\in V$, orients $G$ and annotates joints.
@@ III-B Part-Aware 3D Representations
We generate part-level 3D meshes with a part-aware 3D foundation model [3] from the input RGB images or textual description, and discard meshes with too few vertices or a degenerate spatial spread. For each prospective parent part, we train a continuous SDF [39, 40, 41] $f_{\theta}:\mathbb{R}^{3}\to\mathbb{R}$ on (i) noisy surface points, (ii) near-surface offsets, and (iii) far samples in a bounding AABB.
@@ III-C Kinematic Topology Inference
We choose the base link $b$ as any node in $G$ with the highest undirected degree. Starting from $b$, we run BFS on $G$ as a warm start for the MCTS search. During BFS, when a new neighbour $v$ is first reached from an already visited node $u$, we define $u$ as the parent and $v$ as the child, orient the edge as $u\!\to\!v$, and set its origin $o_{uv}\leftarrow c_{v}-c_{u}$. Any edge whose addition would create a cycle is not inserted into $T$ and is recorded as broken. If $G$ is disconnected, each remaining component is attached to the current tree by adding a virtual edge from its nearest neighbor, provided the Euclidean distance is at most $d_{\max}$.
@@ III-C1 State, actions, constraints
Figure 1: Pipeline of Kinematify for recovering articulated robots from a single RGB image. Step 1: A 3D foundation model generates a segmented mesh of the robot. Step 2: A contact graph is constructed over mesh parts, capturing candidate relations between components. Step 3: Infer the kinematic tree using MCTS, resolving ambiguous connections by leveraging structural priors such as hierarchy and symmetry. Step 4: Refine joint parameters using the DW-CAVL optimization approach while preserving near-contact geometry. Bottom row: Examples of inferred revolute joints with optimized axes.
@@ III-C3 Reward
$R_{\mathrm{struct}}$
This term penalizes irregular depths, excessive out-degree, and long trees:
$R_{\mathrm{struct}}=\frac{1}{\overline{d^{2}}+\overline{(\deg^{+}\!-\!k)^{2}}+\lambda\,|E(\tilde{T})|}$, (1)
where $k$ is the preferred out-degree and $\lambda>0$ is a hyperparameter.
$R_{\mathrm{static}}$
@@
where $\mathrm{ch}(i)$ is the set of children of $i$ in $\tilde{T}$. Let $\hat{\mathbf{z}}^{-}=[0,\,0,\,-1]^{\top}$ denote downward unit gravity and $g>0$ the gravitational constant. The total gravitational torque is
@@ III-D2 DW-CAVL objective
The consistency term penalizes separation near contact after motion,
@@ IV-D Ablation study
Figure 5: Kinematify generates URDFs for both the Fetch robot and the drawer, enabling a demonstration of the robot opening the drawer in Isaac Sim and transfer to real, with the same models usable for online planning with MoveIt.
@@ IV-E Real-world robot manipulation
V Discussion and Limitations
This work addresses the open-vocabulary synthesis of articulated objects from static inputs. However, accurate part segmentation and a mechanically faithful contact graph remain assumptions. Spurious seams or missed contacts can mislead the search or leave subassemblies unattached. A practical extension is to estimate contact reliability and jointly refine segmentation and contacts. We also find that decorative geometry near interfaces can introduce false contacts that bias the search. In such cases, reweighting the MCTS objective to emphasize the symmetry prior, i.e., increasing the weight on $R_{\mathrm{sym}}$ relative to $R_{\mathrm{contact}}$, helps suppress attachments to decorative elements and favors globally coherent parents within symmetric clusters. An adaptive schedule that upweights $R_{\mathrm{sym}}$ when cluster cardinality is high or contact variance is large is a simple and effective mitigation. Learning these weights from validation performance is a complementary direction.
@@ VI Conclusion
We presented Kinematify, an automated pipeline that synthesizes articulated object and robot descriptions directly from RGB images or textual descriptions. Across everyday objects and robot platforms, Kinematify improves joint accuracy and kinematic tree fidelity over prior work. We view Kinematify as a step toward open-vocabulary synthesis of high-DoF articulated structures.